A family invested $10,000 at 8% annual interest. How much additional money must be invested at 9% to obtain a return of 8.2% on the total amount invested?
Answer by mananth(16949)

percent Principal $
8.00 10000
9.00 x
8.20 10000+x
8*10000 +9x=8.2* (10000+x )
80000+9x =8.2*10000+8.2x
80000+9x =82000+8.2x
9x-8.2x= 82000-80000
0.8x=2000
/0.8
x=2500 @ 9.00%